Determination of charge/mass ratio of electron using narrow fine beam tube experiment objective


1
Expert's answer
2022-03-13T18:48:07-0400

The change in kinetic energy of electron when it moves through potential difference "V" is given by

"\frac{mv^2}{2}=eV"

Hence, the speed of electron when it enters in magnetic field

"v=\sqrt{2eV/m}"

The electron in magnetic field moves along a circular path of radius

"r=\frac{mv}{eB}=\frac{m\sqrt{2eV/m}}{eB}=\sqrt{(2mV)/(eB^2)}"

Finally

"\frac{e}{m}=\frac{2V}{r^2B^2}"